1.The water at a local beach has an average depth of 1 metre at low tide. The average depth of the water at high tide is 8 m. If one cycle takes 12 hours:
a.Determine the equation of this periodic function using cosine as the base function where 0 time is the beginning of high tide.
b.What is the depth of the water at 2 am?
c.Many people dive into the beach from the nearby dock. If the water must be at least 3 m deep to dive safely, between what daylight hours should people dive?

Answer by stanbon(75887)

The water at a local beach has an average depth of 1 metre at low tide. The average depth of the water at high tide is 8 m. If one cycle takes 12 hours:
a.Determine the equation of this periodic function using cosine as the base function where 0 time is the beginning of high tide.
y = acos(bx+c) = d
d = 1
amplitude = (7/2) ft.
period = 360/b = 12
b = 30
shift = c/b = 3 implies c = -90
d = 9/2 ft
---
Equation:
f(x) = (7/2)cos(30x - 90) + (9/2)
b.What is the depth of the water at 2 am?
f(2) = (7/2)cos(30*2-90) + (9/2) = 7.53 ft
----------------------------------------------------
c.Many people dive into the beach from the nearby dock. If the water must be at least 3 m deep to dive safely, between what daylight hours should people dive?
I'll leave that to you.
=================================
Cheers,
Stan H.